# A model-based assessment of social isolation practices for COVID-19 outbreak response in residential care facilities

### v1

@@ IV Discussion
Few previous modelling studies approach this subject. Most recent studies focus on the role of vaccination and infection surveillance [26, 3, 9, 27], with some examining case isolation [28], cohorting [29], and furlough of staff [7]). However, our results contrast with one other study of which we are aware. The results reported by Love et al. show substantial benefits from reductions in contact rates between residents, even in combination with case isolation and other measures [23]. However, their model assumes very large differences in the level of contact between residents with and without social distancing (respectively: 2 vs. 50 potentially infectious contacts per resident, per day). The unmitigated contact rate (50 per day) is much higher than those reported by Vilches et al. (6.8 contacts per day) for Canadian care facilities [6], and Smith et al. (5 per day) for a French facility [3], both of which are closer to the rate assumed in our implementation (3 per day). Therefore, while their conclusions may be valid for scenarios with very high baseline contact rates, these rates are not realistic for residential aged care facilities.
@@
IV.1 Limitations
While our model is realistic in some respects, in other ways the level of abstraction it applies leaves room for improvement. One significant limitation arises from the way in which the model simulates furlough of staff members with confirmed infection. The mechanism of case isolation for staff members is to remove them from the facility workforce for a 7-day period. While this response and its duration are realistic, we do not simulate any re-distribution of labour (i.e., re-assignment of rooms to the remaining staff members), an omission which could produce unrealistic reductions in network density and cause the model to over-estimate the effectiveness of furloughing staff with confirmed infection. To realistically capture labour re-distribution patterns (i.e., “surge rostering”) additional consultation is required because specific industry standards are not available.
Regarding the social interface of the RACF with the outside community, our model simulates only a single importation event. We do not explicitly simulate visitors to the facility, or the transmission patterns associated with staff members outside of the facility. Simulating such factors is beyond the scope of this study, but should be considered for future work focused on optimising the criteria for outbreak declaration so to avoid false alarms, or the risks produced by RACF staff working across multiple facilities. Though these topics have been studied by others (e.g., [8]), understanding the role of staffing practices in outbreak prevention and mitigation is applicable to other healthcare sectors as well and many challenges remain unaddressed [35, 36].
@@ IV.1 Limitations
V Conclusion
@@
VI Code Availability
@@
VII Author Contributions
@@
VIII Competing Interests
This work was funded by the Australian Department of Health and Aged Care.
IX Acknowledgements
@@
S1.1 Facility Characteristics
@@
For the results reported in this study, we simulated a single facility with the following values for these parameters:
$n$ $k$ $p_{med}$ $p_{needs}$ $p_{shared}$
88 121 0.3 0.0 0.0
Table S1: Characteristics of the facility used for the simulation results reported in this study.
The number of rooms is determined by the number of residents and the proportion of shared rooms, and assumes that all shared rooms are double-occupancy. Assignment of staff to a work roster follows the data on full-time-equivalent (FTE) and number of sector employees described in the 2020 census of Australian aged care facilities [37]. Rooms are allocated to staff members based on their role (medical or general staff) and are allocated to ensure all rooms are serviced while distributing the number of room assignments as evenly as possible among the workers available on each day of the roster period (1-week). More details are provided below:
@@
S1.2 within-host model of disease progression
The within host model of disease transmission features infectiousness that increases from the time of infection, peaking closely before symptom onset, and then declining until recovery. Crucially, we assume an over-dispersed distribution of peak infectiousness values which corresponds to a negative binomial secondary case distribution, if a sufficiently large number of susceptible contacts exist
@@
where $t_{r}$ is the time between symptom onset and recovery, which is drawn uniformly at random from the range [5d, 10d], to capture a plausible range for the duration of infectiousness after symptom onset.
In the individual-level model of infectiousness described above, the incubation period plays a central role in determining the dynamics of infectiousness. For each individual, these are drawn at random from a lognormal distribution with log-mean $\mu_{inc}=1.62$ and $\sigma=0.418$ (for a mean incubation period of 5.5 days).
To produce the over-dispersed secondary case distribution shown in Figure 2, the individual-level parameter $\beta_{max}$ is drawn from a Gamma distribution with shape parameter $k_{\text{shape}}=0.1$ and scale parameter $s=\kappa/k_{\text{shape}}$, where the mean $\kappa$ globally scales the transmission rate and reproductive ratio.
@@
S1.3 calibration of the reproductive ratio
To calibrate the basic reproductive ratio, we systematically varied the global transmission scalar $\kappa$. For each value of $\kappa$, we simulated an ensemble of 1000 index cases and recorded the number of secondary cases produced for each instance. The mean of each ensemble provided an estimate of the reproductive ratio for the corresponding value of $\kappa$. Figures describing the calibration of $R_{0}(\kappa)$ as well as final size statistics for reach value of $\kappa$ are shown in the Supporting Information Figure S1 and Figure S2. Based on this calibration we selected $\kappa=0.2$ to produce a reproductive ratio $R_{0}\approx 2.4$. To ensure index cases were selected commensurate with their relative likelihood of being infected during an outbreak, index case selection was weighted based on the sum of all contact edge weights associated with each agent (see below for more details about the contact network specifications).
S1.4 test sensitivity
The time-dependent test sensitivity function is derived from a study that performed prospective serial testing of healthcare workers [38] and follows a similar trajectory to infectiousness, with sensitivity peaking just before symptom onset and a gradual decline afterwards. Test sensitivity of rapid antigen tests is assumed to peak at approx. 83% on average (though this varies from person to person). Extensive details about the specifics of the test sensitivity model used here can be found in our previous work [19]. We assume tests are 100% specific (no false positivies).
S1.5 network model of disease transmission
@@
S1.5.1 Component (i): structured contact network
Component (i) of the transmission network is a dynamic weighted multigraph (multiple edges may exist between two nodes) that includes both staff members and residents as nodes. The edges in the network connect staff members who are assigned to service the same room on the same day, residents to their roommates, and residents to the staff members who service their rooms. Figure 1 provides a schematic description of how this part of the network is generated, based on the assignments of residents and staff to rooms. The network topology of component (i) is dynamic, changing for each day of the simulation based on which staff are assigned to which rooms. Furlough of staff members who test positive also changes the network topology by removing all connections with furloughed staff members. These connections are re-instated after the 7-day furlough period. The edges in component (i) are weighted, to account for the relative likelihood that a given edge is sampled during the transmission simulations (described below). Specifically, edge weights depend pairwise on the types of agents they connect as shown in Table S2. For staff members who are assigned more than one of the same rooms, an edge exists between them for each room they share.
@@
S1.5.2 Component (ii): random contacts between residents
@@
S1.5.3 Contact sampling
In each discrete simulation timestep (of duration $\Delta t=0.1d$), pairwise transmission dynamics are evaluated between infected individuals, and sampled subsets of their potential contacts. This sampling process is split into components (i) and (ii) as described above. Here, we describe the details of our contact sampling algorithm for components (i) and (ii). We note that the network topology of component (i) is a function both of the day of the staff roster cycle and of the set of staff who are furloughed and not present at the facility. In the description below, we omit notation associated with these two aspects of network structure dynamics, in order to clearly denote the process used to sample infectious contacts from the set allowed by network structure at any time $t$.
@@
The number of potentially infectious contacts from component (i) $c_{\text{(i)}}(t)$ sampled for a given timestep is then drawn from a Poisson distribution:
@@
S1.6 Pairwise disease transmission
@@
where $\beta_{a}$ is the force of infection produced by the infected agent $a$, $\tau$ is the time since individual $a$ was initially infected, $\eta_{ab}$ is a reduction factor for PPE effects (which depend on whether individuals $a$ and $b$ are residents or staff members), and $S_{b}$ takes a value of 00 if individual $b$ is infected or recovered and $1$ if they are susceptible. Note that the role of the timestep $\Delta t$ is not included in Equation because it is accounted for in the contact rates described in Equations S4 and S6. The transmission rate $\beta_{a}$ changes as a function of time since infection, as described by our within-host model of disease transmission which accounts for dispersion of infectiousness among infected individuals, and the timing of transmission as described in detail in our previous work [19].
S1.7 implementation of interventions and PPE deployment
The assumptions listed below with respect to screening for cases and response to outbreaks were informed by the advice detailed in the Australian Government Department of Health document entitled ”COVID-19 Outbreaks in Residential Care Facilities, Communicable Disease Network Australia, National Guidelines for the Prevention, Control and Public Health Management of COVID-19 Outbreaks in Residential Care Facilities” dated February 15th, 2022 [18]. [Note: the Australian Government recommendations have been modified since this study was conducted, the version of the guidelines we used to inform our model is included in supporting material to ensure continued access.]
@@
    Staff screening (no active outbreak). Staff are tested periodically, with no more than two tests scheduled per roster period (weekly) and a minimum of 3 days between subsequent tests. This means full-time workers are scheduled for two tests per week and part time workers are scheduled for one test per week.
@@
    Resident screening (no active outbreak). Residents are not subject to scheduled testing, but have a baseline probability of being tested on any given day.
@@
    type of interaction % reduction in force of infection
    resident $\rightarrow$ resident 20%
    worker $\rightarrow$ worker 50%
    worker $\rightarrow$ resident 90%
@@
    Throughout, we assume a 90% reduction in background contact frequency for residents who are isolated after testing positive. This reflects the imperfect capacity to isolate a resident to their room at all times. [We note here that contact rates with staff and with roommates are unaffected by resident case isolation].
@@
    Furlough of staff members with confirmed infection continues for 7 days after the day of detection. For the 7 day period, a furloughed staff member is removed from the model (their connections are removed from the transmission network). We do not simulate re-distribution of room assignments upon furlough of staff, rather, we assume that all network connections not involving furloughed staff members remain unchanged.
@@
    Declaring an active outbreak: For a facility to declare an active outbreak, one of two criteria must be met: (i) two new resident cases within 5 days or (ii) 5 new cases in staff within 7 days as per the Australian Government guidelines [18].
@@
    Declaring an active outbreak over: If no new resident cases have been detected within 7 days, the outbreak is considered to be over, and infection control measures cease.
S1.8 vaccine-derived protection

### v2

@@ V Discussion
Few previous modelling studies examine general isolation policies. Most recent studies focus on the role of vaccination and infection surveillance [36, 5, 11, 37], with some examining case isolation [38], cohorting [39], and furlough of staff [9]). However, our results contrast with one other study of which we are aware. The results reported by Love et al. show substantial benefits from reductions in contact rates between residents, even in combination with case isolation and other measures [30]. However, their model assumes very large differences in the level of contact between residents with and without social distancing (respectively: 2 vs. 50 potentially infectious contacts per resident, per day). The unmitigated contact rate (50 per day) is much higher than those reported by Vilches et al. (6.8 contacts per day) for Canadian care facilities [8], and Smith et al. (five per day) for a French facility [5], both of which are closer to the rate assumed in our implementation (three per day). Therefore, while their conclusions may be valid for scenarios with very high baseline contact rates, these rates are not realistic for residential aged care facilities.
@@
V.1 Limitations
While our model is realistic in its representation of key factors such as the facility structure, staffing schedules, within-host disease dynamics, test sensitivity, and outbreak response, in other ways the level of abstraction it applies introduces some limitations. One of these is the way in which the model simulates furlough of staff members with confirmed infection. The mechanism of case isolation for staff members is to remove them from the facility workforce for a 7-day period. While this response and its duration are realistic, we do not simulate re-assignment of rooms to the remaining staff members, an omission which could produce unrealistic reductions in network density and cause the model to over-estimate the effectiveness of furloughing staff with confirmed infection. To realistically capture labour re-distribution patterns (i.e., “surge rostering”) additional consultation is required because specific industry standards are not available. This limitation allows residents to be completely isolated due to the furlough of all staff servicing their room. We confirmed that such scenarios arise to a negligible extent (and do not occur at all in more than 90% of simulations, see Supporting Information Figure S1).
Regarding the social interface of the RACF with the outside community, our model simulates only a single importation event. We do not explicitly simulate visitors to the facility, or the transmission patterns associated with staff members outside of the facility. Our model therefore implicitly assumes that visitors play a negligible role in transmission during active outbreaks, and that there is a low likelihood of multiple importations occurring prior to outbreak declaration. In scenarios with high community prevalence, multiple importation events from visitors could feasibly occur and could lead to larger numbers of presymptomatic or asymptomatic infections at the time of outbreak declaration. However, multiple importations would not alter the reproductive ratio, so we do not expect our main results regarding the marginal impact of general isolation after outbreak detection to be sensitive to the number of initial infections. Looking to future work, we note that restrictions on visitation were among the most heavily criticised policies applied to aged care facilities during COVID-19 and this subject deserves further attention focused on balancing the well-being of residents and their families with the need to prevent importation of infectious diseases. In general, while simulating the community external to the facility is beyond the scope of this study, it should be considered for future work on topics such as optimising the criteria for outbreak declaration so to avoid false alarms, or the risks produced by RACF staff working across multiple facilities. Though these topics have been studied by others (e.g., [10]), understanding the role of staffing practices in outbreak prevention and mitigation is applicable to other healthcare sectors as well and many challenges remain unaddressed [45, 46].
@@ V.1 Limitations
Finally, our choice to assume a uniform asymptomatic fraction of 33% neglects the potential for SARS-CoV-2 infection to present higher clinical fractions in older age groups. Assuming higher symptomatic proportions in the resident population due to their older age could reduce the marginal impact of asymptomatic screening. However, even with a low asymptomatic fraction, screening would still play an important role in early case identification during incubation, limiting the sensitivity of our findings to this simplifying assumption.
VI Conclusion
@@
VII List of Abbreviations
RACF - residential aged care facility; PPE - personal protective equipment; NPIs - nonpharmaceutical interventions; FTE - full-time equivalent;
VIII Declarations
VIII.1 Ethics approval and consent to participate
Not applicable
VIII.2 Consent for publication
Not applicable
VIII.3 Availability of data and materials
@@
VIII.4 Competing Interests
Cameron Zachreson: Reports no competing interests.
Ruarai Tobin: Reports no competing interests.
Camelia Walker: Reports no competing interests.
Eamon Conway: Reports no competing interests.
Freya M Shearer: Reports payments made to the University of Melbourne by DMTC Limited for supporting the development of an epidemic decision support system.
Jodie McVernon: Reports payments to the University of Melbourne by the World Health Organisation and payments to the University of Melbourne by the Australian Department of Foreign Affairs and Trade.
Nicholas Geard: Reports no competing interests
VIII.5 Funding
This work was funded by the Australian Department of Health and Aged Care and by the Australian Government’s Medical Research Future Fund (MRF2017355).
VIII.6 Author Contributions
@@
VIII.7 Acknowledgements
@@
S1.1 Overview
The model used in this work consists of three main components: 1) a structured population contact network model based on the characteristics of an RACF environment, 2) a model of COVID-19 within-host disease progression and transmission between infected and susceptible individuals, and 3) a model of case detection and outbreak response. Pesudocode describing the overall simulation algorithm is shown in Algorithm S1 and a more detailed description of the transmission algorithm providing key details of the transmission model implementation is provided in Algorithm S2.
A list of key model parameters is provided in Tables S1, S2, and S3. The method descriptions provided in the following sections provide details of model implementation. The method description below does not contain mathematical details of the dynamic model of test sensitivity, which is described comprehensively in our previous work [21].
Algorithm S1 Overview of the model algorithm.
population $\leftarrow$ generate population(facility characteristics)
network $\leftarrow$ initialise contact network(population)
scenario $\leftarrow$ set scenario configuration(input parameters)
configuration $\leftarrow$ (population, network, scenario)
while outbreaks $<$ 1000 do
     initial conditions $\leftarrow$ initialise dynamics(configuration)
     (outbreak declared, stats) $\leftarrow$ transmission(initial conditions, configuration)
     if outbreak declared then
         outbreaks $\leftarrow$ outbreaks + 1
         push stats to ensemble output
     end if
end while
return summary statistics for all outbreaks
Algorithm S2 Outbreak simulation algorithm
function transmission(initial conditions, configuration)
     outbreak declared $\leftarrow$ false
     termination flag $\leftarrow$ false
     while termination flag is false do
         iterate transmission dynamics:
         $t\leftarrow t+\Delta t$
         update active infections(population) $\triangleright$ progress within-host model
         update environment:
         if new day then
              update structured contact network(day of week, network)
              test staff and residents(outbreak declared, scenario, population)
              update case isolation(outbreak declared, scenario, population)
              outbreak response:
              if outbreak declaration conditions are met then
                  outbreak declared $\leftarrow$ true
                  outbreak response(scenario) $\triangleright$ alters testing, PPE, distancing
              end if
              if termination conditions are met then
                  termination flag $\leftarrow$ true
              end if
         end if
         contact sampling:
         initialise potential transmission pairs: $E_{I}\leftarrow[~{}]$
         initialise aggregate weight of infectious edges: $w_{I}\leftarrow 0$
         for $a$in infected agents do
              push edges $e_{a}$ to $E_{I}$triangleright compile list of infectious edges
              $w_{I}\leftarrow w_{I}+\sum\limits_{e\in e_{a}}w_{e}$triangleright sum weights of infectious edges
         end for
         sample infectious edges: $E_{T}\subset E_{I}:p(e\in E_{T})\propto w_{e}/w_{I}$
         push background contacts to $E_{T}$triangleright random contacts between residents
         pariwise transmission:
         for $e\in E_{T}$ do
              evaluate pairwise transmission probability
              if transmission successful then infect susceptible contact
              end if
         end for
     end while
     return outbreak declared, summary stats
end function
Parameters: RACF facility model
Symbol Description Value
$n$ Number of residents 88
$k$ Number of facility staff 121
$p_{med}$ Prop. of medically-trained staff 0.3
$p_{needs}$ Prop. of high-needs residents 0
$p_{shared}$ Prop. of double-occupancy rooms 0
$p_{\text{min}}$ minimum fraction of staff present each day in baseline roster 0.2
$p_{5}$ prop. of staff working 5 days per week 0.26
$p_{3}$ prop. of staff working 3 days per week 0.36
$p_{2}$ prop. of staff working 2 days per week 0.38
$k_{gen}$ number of general staff assigned to each room per day 4
$k_{med}$ number of medical staff assigned to each room per day 1
Structured contact network
$c$ mean number of contacts per resident, per day 3.0
$\lambda_{\text{(i)}}$ mean total contact rate (contacts per day) $nc$
$w_{e}$ sampling weight for each edge see Table S4
Unstructured contact network
$\lambda_{\text{(ii)}}$ mean number of background contacts per resident per day (baseline, no active outbreak) 3.0
Table S1: Summary of key model parameters related to population structure. For additional details regarding the facility model and contact networks see the corresponding sections below.
Parameters: COVID-19 transmission model
Symbol Description Value
$\kappa$ global transmission scalar 0.2
$R_{0}$ mean reproductive ratio $\approx 2.4$
$\delta$ dispersion parameter of secondary case distribution 0.1
$\beta_{max}$ random variable determining an individual’s peak infectiousness Gamma (shape = $\delta$, scale = $\kappa/\delta$)
$T_{\text{inc}}$ mean incubation period 5.5 days
$t_{\text{inc}}$ random variable determining an individual’s incubation period Lognormal ($\mu=1.62$, $\sigma=0.418$)
$T_{\text{rec}}$ mean recovery period 7.5 days
$t_{\text{rec}}$ an individual’s recovery period Uniform(5, 10)
$p_{\text{asymp}}$ mean proportion of asymptomatic infections 0.33
Table S2: Summary of key model parameters. For additional details on the COVID-19 transmission model, see the relevant section below and our previous work [21].
Parameters: Outbreak detection and response
Symbol Description Value
$\rho_{max}$ peak test sensitivity (see [21] and [47] for further details) $\approx 0.83$
$\Delta^{\text{staff}}_{\text{min}}$ minimum time (in days) between asymptomatic tests for staff members (no active outbreak) 3
$p^{\text{resident}}_{\text{test}}$ probability per day of an asymptomatic resident testing for infection 0 or 1 (depends on scenario)
$p^{\text{symptoms}}_{\text{test}}$ probability of testing if symptomatic (note: staff members are only tested if present) 1.0
$t_{\text{furlough}}$ duration of furlough period for staff who test positive 7 days
$t_{\text{iso}}$ duration of case isolation period for residents who test positive 7 days
$p_{\text{iso}}$ case isolation efficacy 0.9
$L$ general isolation efficacy 0.9, 0.5, or 0 (depends on scenario)
$\eta_{ab}$ efficacy of infection control (e.g., PPE) 0.2, 0.5, or 0.9 (depends on context see Table S5)
Table S3: Summary of key model parameters. For additional details regarding the outbreak response model (including screening strategies, criteria for declaring outbreaks and criteria for declaring outbreaks to be over) see the corresponding sections below.
S1.2 Facility Characteristics
@@
For the results reported in this study, we simulated a single facility with the values for these parameters listed in Table S1. The number of rooms is determined by the number of residents and the proportion of shared rooms, and assumes that all shared rooms are double-occupancy (Note that in this work we assume all rooms are single-occupancy).
Assignment of staff to a work roster follows the data on full-time-equivalent (FTE) and number of sector employees described in the 2020 census of Australian aged care facilities [3]. Rooms are allocated to staff members based on their role (medical or general staff) and are allocated to ensure all rooms are serviced while distributing the number of room assignments as evenly as possible among the workers available on each day of the roster period (1-week). More details are provided below:
@@
S1.3 within-host model of disease progression
The within host model of disease transmission features infectiousness that increases from the time of infection, peaking closely before symptom onset, and then declining until recovery. Crucially, we assume an over-dispersed distribution of peak infectiousness values which corresponds to a negative binomial secondary case distribution, if a sufficiently large number of susceptible contacts exist.
@@
where $t_{r}$ is the time between symptom onset and recovery, which is drawn uniformly at random from the range [5d, 10d], to capture a plausible range for the duration of infectiousness after symptom onset with a mean of $T_{\text{rec}}=7.5$ days.
In the individual-level model of infectiousness described above, the incubation period plays a central role in determining the dynamics of infectiousness. For each individual, these are drawn at random from a lognormal distribution with log-mean $\mu_{inc}=1.62$ and $\sigma=0.418$ (for a mean incubation period of $T_{\text{inc}}=5.5$ days). For asymptomatic infections, the time-dependent infectiousness trajectory is computed identically as for those which do express symptoms. The only difference is that symptoms cannot be observed after incubation and symptomatic screening for case detection is not conducted. We simulate a mean asymptomatic fraction of $33\%$ using a probabilistic implementation: at the moment of infection, an individual will be labeled as asymptomatic with a probability of $p_{\text{asymp}}=0.33$. This asymptomatic fraction is in line with global estimates but we do not account for heterogeneity by age [23, 24, 25, 26].
To produce the over-dispersed secondary case distribution shown in Figure 2, the individual-level parameter $\beta_{max}$ is drawn from a Gamma distribution with shape parameter $k_{\text{shape}}=0.1$ and scale parameter $s=\kappa/k_{\text{shape}}$, where the mean $\kappa$ globally scales the transmission rate and reproductive ratio. This implementation produces a negative-binomial secondary case distribution with dispersion $\delta=0.1$, equivalent to the shape parameter of the Gamma distribution for $\beta_{max}$.
@@
S1.4 calibration of the reproductive ratio
To calibrate the basic reproductive ratio, we systematically varied the global transmission scalar $\kappa$. For each value of $\kappa$, we simulated an ensemble of 1000 index cases and recorded the number of secondary cases produced for each instance. The mean of each ensemble provided an estimate of the reproductive ratio for the corresponding value of $\kappa$. Figures describing the calibration of $R_{0}(\kappa)$ as well as final size statistics for reach value of $\kappa$ are shown in Figure S2 and Figure S3. Based on this calibration we selected $\kappa=0.2$ to produce a reproductive ratio $R_{0}\approx 2.4$. To ensure index cases were selected commensurate with their relative likelihood of being infected during an outbreak, index case selection was weighted based on the sum of all contact edge weights associated with each agent (see below for more details about the contact network specifications).
S1.5 test sensitivity
The time-dependent test sensitivity function is derived from a study that performed prospective serial testing of healthcare workers [47] and follows a similar trajectory to infectiousness, with sensitivity peaking just before symptom onset and a gradual decline afterwards. Test sensitivity of rapid antigen tests is assumed to peak at approx. 83% on average (though this varies from person to person). Extensive details about the specifics of the test sensitivity model used here can be found in our previous work [21]. We assume tests are 100% specific (no false positives).
S1.6 network model of disease transmission
@@
S1.6.1 component (i): structured contact network
Component (i) of the transmission network is a dynamic weighted multigraph (multiple edges may exist between two nodes) that includes both staff members and residents as nodes. The edges in the network connect staff members who are assigned to service the same room on the same day, residents to their roommates, and residents to the staff members who service their rooms. Figure 1 provides a schematic description of how this part of the network is generated, based on the assignments of residents and staff to rooms. The network topology of component (i) is dynamic, changing for each day of the simulation based on which staff are assigned to which rooms. Furlough of staff members who test positive also changes the network topology by removing all connections with furloughed staff members. These connections are reinstated after the 7-day furlough period. The edges in component (i) are weighted, to account for the relative likelihood that a given edge is sampled during the transmission simulations (described below). Specifically, edge weights depend pairwise on the types of agents they connect as shown in Table S4. For staff members who are assigned more than one of the same rooms, an edge exists between them for each room they share.
@@
S1.6.2 component (ii): random contacts between residents
@@
S1.6.3 contact sampling
In each discrete simulation timestep (of duration $\Delta t=0.1$ day), pairwise transmission dynamics are evaluated between infected individuals, and sampled subsets of their potential contacts. This sampling process is split into components (i) and (ii) as described above. Here, we describe the details of our contact sampling algorithm for components (i) and (ii). We note that the network topology of component (i) is a function both of the day of the staff roster cycle and of the set of staff who are furloughed and not present at the facility. In the description below, we omit notation associated with these two aspects of network structure dynamics, in order to clearly denote the process used to sample infectious contacts from the set allowed by network structure at any time $t$.
@@
i.e., the expected total number of contacts during the timestep multiplied by the proportion of infected edges. The number of potentially infectious contacts from component (i) $c_{\text{(i)}}(t)$ sampled for a given timestep is then drawn from a Poisson distribution:
@@
S1.7 pairwise disease transmission
@@
where $\beta_{a}$ is the force of infection produced by the infected agent $a$, $\tau$ is the time since individual $a$ was initially infected, $\eta_{ab}$ is a reduction factor for PPE effects (which depend on whether individuals $a$ and $b$ are residents or staff members), and $S_{b}$ takes a value of 00 if individual $b$ is infected or recovered and $1$ if they are susceptible. Note that the role of the timestep $\Delta t$ is not included in Equation S10 because it is accounted for in the contact rates described in Equations S4 and S6. The transmission rate $\beta_{a}$ changes as a function of time since infection, as described by our within-host model of disease transmission which accounts for dispersion of infectiousness among infected individuals, and the timing of transmission as described in detail in our previous work [21].
S1.8 implementation of interventions and PPE deployment
The assumptions listed below with respect to screening for cases and response to outbreaks were informed by the advice detailed in the Australian Government Department of Health document entitled “COVID-19 Outbreaks in Residential Care Facilities, Communicable Disease Network Australia, National Guidelines for the Prevention, Control and Public Health Management of COVID-19 Outbreaks in Residential Care Facilities” dated February 15th, 2022 [20]. [Note: the Australian Government recommendations have been modified since this study was conducted, the version of the guidelines we used to inform our model is included as Supporting Material to ensure continued access.]
@@
    Staff screening (no active outbreak). Staff are tested periodically, with no more than two tests scheduled per roster period (weekly) and a minimum of $\Delta^{\text{staff}}_{\text{min}}=3$ days between subsequent tests. This means full-time workers are scheduled for two tests per week and part time workers are scheduled for one test per week.
@@
    Resident screening (no active outbreak). Residents are not subject to scheduled testing, but have a baseline probability $p^{\text{resident}}_{\text{test}}$ of being tested on any given day which varies depending on the screening scenario.
@@
    type of interaction reduction in force of infection ($\eta_{ab}$)
    resident $\rightarrow$ resident 0.2
    worker $\rightarrow$ worker 0.5
    worker $\rightarrow$ resident 0.9
    Table S5: Efficacy of infection control measures (e.g., PPE), which depends on the types of individuals interacting ($\eta_{ab}$).
@@
    Throughout, we assume a 90% reduction in background contact frequency for residents who are isolated after testing positive ($p_{\text{iso}}=0.9$). This reflects the imperfect capacity to isolate a resident to their room at all times. [We note here that contact rates with staff and with roommates are unaffected by resident case isolation]. Case isolation lasts for a fixed period of $t_{\text{iso}}=7$ days.
@@
    Furlough of staff members with confirmed infection continues for $t_{\text{furlough}}=7$ days after the day of detection. For the seven day period, a furloughed staff member is removed from the model (their connections are removed from the transmission network). We do not simulate re-distribution of room assignments upon furlough of staff, rather, we assume that all network connections not involving furloughed staff members remain unchanged. This produces a small probability that residents will become completely isolated from all staff members servicing their rooms (an unrealistic condition). For quality control, we quantified this probability for the scenario producing the highest FTE deficit (this occurs when asymptomatic screening is implemented upon outbreak declaration, with general isolation level set to zero, see Table 1). The analysis summarised in Figure S1 demonstrates negligible levels: for a sample of 1000 simulated outbreaks, 90.1% showed zero resident-days of isolation due to staff furlough, 5.5% of simulated outbreaks produced one resident-day of isolation due to staff furlough, 3.5% produced two to five resident-days, and 0.2% produced six to nine resident-days of isolation due to staff furlough.
@@
    Declaring an active outbreak: For a facility to declare an active outbreak, one of two criteria must be met: (i) two new resident cases within 5 days or (ii) five new cases in staff within seven days as per the Australian Government guidelines [20].
@@
    Declaring an active outbreak over: If no new resident cases have been detected within seven days, the outbreak is considered to be over, and infection control measures cease.
Figure S1: Frequency of resident-days of isolation due to furlough of staff members over 1000 outbreaks. For these simulations, the condition producing the maximum FTE deficit was selected (asymptomatic screening during outbreaks, with 0% general isolation level, see Table 1). Over the ensemble of 1000 simulated outbreaks, 90.1% produced zero, 5.5% produced one, 3.5% produced two to five, and 0.2% produced six to nine resident-days of isolation due to staff furlough.
S1.9 vaccine-derived protection
@@
S4 Sensitivity to background contact rate $\lambda_{(\text{ii})}$ and case isolation compliance
Our main results on the marginal impact of general isolation conditions are sensitive to the rate of contact between residents (reduced through general isolation) relative to the rate of contact in the structured network (unaffected by general isolation). To investigate this sensitivity, we fix the number of structured contacts per resident fixed at $c=3.0$, and compute outbreak final size while systematically varying the unstructured contact rate per resident ($\lambda_{(\text{ii})}$). For fair comparisons between scenarios, we fix the reproductive ratio $R_{0}$ by jointly varying $\kappa$ to control the transmission rate for each value of $\lambda_{(\text{ii})}$. The results in Figure S5 demonstrate that our primary results hold when the ratio $\lambda_{(\text{ii})}/c$ is near one. Even for larger values of $\lambda_{(\text{ii})}/c$, the marginal impact of general isolation decreases substantially as the reproductive ratio decreases (Figure S5).
The sensitivity of our main results to the level of compliance with case isolation is shown in Figure S6, which demonstrates that marginal impacts of general isolation are not sensitive to assumptions regarding case isolation compliance levels. The robustness of this result depends on the realistic assumption that general isolation compliance is bounded from above (cannot be greater than) compliance with targeted case isolation.
Figure S5: Sensitivity analysis of background contact rate on the marginal impact of general isolation. Sample medians (n = 1000 outbreaks) are shown as functions of the ratio of the unstructured contact rate per resident $\lambda_{(\text{ii})}/c$, where $c$ is the rate of contact associated with staff members and is unaffected by general isolation measures for residents. Results are shown for different values of the reproductive ratio $R_{0}$ (a-d), with and without the application of general isolation for outbreak control. Marginal impacts of general isolation increase with $R_{0}$ and $\lambda_{(\text{ii})}$.
Figure S6: Sensitivity analysis of case isolation compliance on the marginal impact of general isolation. Sample medians (n = 1000 outbreaks) are shown as functions of case isolation compliance rates for residents. Results are shown for four different values of $R_{0}$ with and without the application of general isolation for outbreak control. Here, we assume compliance with general isolation is equal to the compliance with case isolation.